OpenStudy (anonymous):

Solve the system of equations by the addition method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. -4x - 3y = 6 -20x - 15y = 30 A. No solution, inconsistent B. (-24, -18) C. (-18, -24) D. Infinitely many solutions, dependent

OpenStudy (anonymous):

Try dividing the second equation by 5 and you can see what will happen. This is a pretty easy problem. You can probably do it by just looking at it.

OpenStudy (anonymous):

I thought it was A math isn't really my easy subject so you would be surprised how hard this is for me lol...

OpenStudy (anonymous):

can you show me in steps like lil steps

OpenStudy (anonymous):

Ok let me help you. You have two equations, \[-4x-3y=6\]\[-20x-15y=30\]Now if you divide the second equation by 5, it means you divide every number by 5, you get:\[-\frac{ 20 }{ 5 }x - \frac{ 15 }{ 5}y = \frac{ 30 }{ 5 }\]and the resulting second equation now is:\[-4x-3x=6\]and it is easy to see that both your first and second equations are the same. If you plot this you can see that they are on top of each other. And just like everything on top of teach other, they touch everywhere therefore you have infinitely many solutions, dependent on an integer n. That is - it depends on you to select a point and you can get the solution for that. Was it easy ;)

OpenStudy (anonymous):

Lol the way you put it was easy x)